CLICK HERE TO REVIEW LEARNING OBJECTIVES QUESTION 5 Not complete Marked out of 16.00 Flag question Two-Stage ABC for Manufacturing Assume Sherwin-Williams Company, a large paint manufacturer, has determined the following activity cost pools and cost driver levels for the latest period Activity Cost Pool Activity Cost Activity Cost Driver Machine setup Material handling Machine operation $1,000,000 870,000 200,000 2,500 setup hours 5,000 material moves 20,000 machine hours The following data are for the production of single batches of two products, Mirlite and Subdue: Mirlite Subdue Gallons produced Direct labor hours Machine hours Direct labor cost Direct materials cost Setup hours Material moves 50,000 30,000 250 250 $11,000 $8,500 $200,000 12 35 400 800 $400,000 Determine the batch and unit costs per gallon of Mirlite and Subdue using ABC. Round cost per gallon to two decimal places. Product Costs Mirlite Subdue Direct materials Direct labor 0 Manufacturing overhead: Machine setups 0 Material handling Machine operation Total cost per batch Gallons produced 0 Cost per gallon Check
Answer #1
Product Cost
Mirlite Subdue
Direct Materials $ 400,000.00 $ 200,000.00
Direct Labor $    11,000.00 $    85,000.00
Manufacturing Overhead
Machine Setups $      6,000.00 $      4,800.00
Material Handling $    10,440.00 $      6,090.00
Machine Operation $      8,000.00 $      2,500.00
Total Cost Per Batch $ 435,440.00 $ 298,390.00
Gallons Produced $    50,000.00 $    30,000.00
Cost Per Gallon $               8.71 $               9.95

Cost Per Activity

Machine Set up= 1,000,000/2500 =400 per Setup Hours

Material Handling= 870,000/5,000 = 174 Per Material Moves

Machine Operation= 200,000/20,000 =10 Per Machine Hours